# Do RAG Systems Suffer From Positional Bias?

### v1

@@ Abstract
Retrieval Augmented Generation enhances LLM accuracy by adding passages retrieved from an external corpus to the LLM prompt. This paper investigates how positional bias—the tendency of LLMs to weight information differently based on its position in the prompt—affects not only the LLM’s capability to capitalize on relevant passages, but also its susceptibility to distracting passages. Through extensive experiments on three benchmarks, we show how state-of-the-art retrieval pipelines, while attempting to retrieve relevant passages, systematically bring highly distracting ones to the top ranks, with over 60% of queries containing at least one highly distracting passage among the top-10 retrieved passages. As a result, the impact of the LLM positional bias, which in controlled settings is often reported as very prominent by related works, is actually marginal in real scenarios since both relevant and distracting passages are, in turn, penalized. Indeed, our findings reveal that sophisticated strategies that attempt to rearrange the passages based on LLM positional preferences do not perform better than random shuffling.
Do RAG Systems Suffer From Positional Bias?
Florin Cuconasu1,2thanks: Work conducted while FC being a research intern at TII.thanks: cuconasu@diag.uniroma1.it, Simone Filice2thanks: filice.simone@gmail.com Guy Horowitz2, Yoelle Maarek2, Fabrizio Silvestri1 1Sapienza University of Rome, 2Technology Innovation Institute
@@ 1 Introduction
Retrieval Augmented Generation (RAG) improves the factual accuracy of LLMs on knowledge-intensive tasks by including in the prompt passages retrieved from an external corpus (Chen et al., 2017; Petroni et al., 2021b; Fan et al., 2024). Because any real retriever is imperfect, RAG systems feed the LLM several top-ranked passages, not just the single best one. That practice raises recall but also inserts distracting passages—text that looks relevant yet lacks the appropriate answer. Recent work shows these distractors can sharply degrade the LLM answer accuracy (Cuconasu et al., 2024; Jin et al., 2025; Yoran et al., 2024).
@@ 2 Related work
Recent work explores the detrimental effect of irrelevant content in the LLM prompt. In the RAG setting, a passage is considered irrelevant if it does not provide useful information for answering the query. Cuconasu et al. (2024) divide irrelevant passages as either random, if they are semantically unrelated to the query, or distracting, if they are related to the query but do not contain the answer. They show that while random passages do not affect answer quality, distracting passages do. Jin et al. (2025) show that irrelevant passages returned by strong retrievers are more detrimental than those obtained by weak retrievers. Amiraz et al. (2025) propose a continuous measure of the distracting effect of irrelevant passages and a fine-tuning approach to enhance LLM robustness, similar to strategies in Lin et al. (2024); Jin et al. (2025); Yoran et al. (2024).
@@ 3 Experimental Setup
Following Zheng et al. (2023); Gu et al. (2025); Rahmani et al. (2024), we evaluate passage relevance and answer quality using the LLM-as-a-judge approach. In the former case, we prompt the LLM to assess the relevance of a passage to a question given the ground truth answer; in the latter, we prompt the LLM to assess whether the generated response semantically matches the reference answer [^3] Exact prompts are provided in Appendix A.3.. We use Claude 3.7 Sonnet via AWS Bedrock as the backbone LLM.
@@
$\text{DE}_{q}(p)=1-p^{\text{LLM}}(\text{NO-RESPONSE}|q,p)\vspace{-0.1cm}$ (1)
For each retrieval pipeline, we compute the distracting effect of the retrieved irrelevant passages and assume DE=0 for relevant passages. Fig. 1c reports the DE of the most distracting passage among the top-$k$ positions (MaxDE), while Fig. 1d reports the mean DE considering the top-$k$ positions (MeanDE). Both metrics are averaged across all queries. The MaxDE curves reach very high values, with Table 4 (Appendix) showing that over 60% of queries contain at least one hard distractor (defined as having a DE score greater than 0.8) in the top-10 results from dense retrievers. The MeanDE curves are initially very low, since most of the top retrieved passages are relevant, then increase as more irrelevant passages appear in the prompt, but soon they decrease again. This suggests that highly distracting passages typically appear in top positions, while low-ranked passages have a DE score close to 0. Finally, retrieval pipelines leading to higher HITS and Precision, e.g., when using BGE, also exhibit higher MaxDE and MeanDE curves, revealing a critical aspect: stronger retrievers increase recall and deliver more harmful distractors, making retrieval a double-edged sword.
@@ 4 Positional Bias in Controlled Settings
For each query, we select the highest-ranked relevant passage obtained by BGE large after reranking. Following Amiraz et al. (2025), we compute the distracting effect for irrelevant passages using Equation 1. We classify passages as “hard distractors” (with DE $>0.8$, as previously defined) and “weak distractors” (with DE $<0.2$). Fig. 2 shows results for Qwen 2.5 7B (results for other models and single datasets are given in Appendix B). Fig. 2a displays the characteristic U-shaped accuracy pattern when rotating a single relevant passage among fixed weak distractors [^4] We use weak distractors instead of general retrieved irrelevant passages to avoid negative effects from hard distractors.. Fig. 2b shows that this positional bias extends to distracting passages, with hard distractors at the beginning or end having significantly higher distracting effect (36-44%) compared to middle slots (28-34%) [^5] We calculate the distracting effect using Equation 1 applied to the entire set of passages rather than a single passage.. This parallel pattern indicates the model favors certain positions regardless of passage relevance.
@@ 5 Positional Bias in Real Scenarios
Results in Tables 2 and 3 show that the impact of the positional bias in real settings is minor: different passage arrangement strategies lead to very similar results that do not significantly differ from the Shuffle baseline [^8] Statistical significance using Wilcoxon test with $p$=0.05., regardless of the LLM or the retrieval pipeline. We argue that these results can be explained by the contrastive effect of relevant and highly distracting passages, which, as observed in Fig. 1, tend to both appear in top retrieved passages: for instance, in the MaxRelevance strategy, the benefit of placing relevant passages in LLM-favored positions is compensated by the unintended tendency to put in the same slots highly distracting passages.
@@ 6 Conclusions
Our work demonstrates that while positional bias exists in current LLMs, its impact is minimal in realistic RAG settings: random ordering of retrieved passages yields statistically equivalent accuracy to more sophisticated reordering strategies. We observed that contemporary retrievers do not merely return some irrelevant passages, they surface passages that degrade answer accuracy in more than 60% of our test questions, turning the retriever itself into a first-order source of error. Thus, attempting to place relevant passages in LLMs’ favorable positions may inadvertently prioritize hard distractors over relevant content, counterbalancing the potential benefits of strategic reordering. These findings suggest that future improvements should focus on retrieval quality and LLM distraction robustness rather than passage positioning.
@@ A.2 Additional Retrieval Results
Figures 3, 4 and 5 report the retrieval results of BM25 and BGE with and without re-ranker (RR) on PopQA, NQ, and TriviaQA, respectively.
@@ B.1 Positional Bias in Controlled Settings
Figures 9, 10, 11 and 12 illustrate the positional bias in controlled settings when rotating either the relevant passage or a hard distractor among weak distractors. The results reveal that each model exhibits its own characteristic positional pattern, confirming findings from Hutter et al. (2025).
@@
Among the LLMs tested, Qwen 2.5 7B demonstrates the most pronounced positional bias (see Fig. 9), while the Llama 3 family appears more resilient to position changes (see Figures 10, 11 and 12). A possible explanation is that these models may have been specifically trained to mitigate the lost-in-the-middle effect. Since this problem has become well-documented in the literature (Liu et al., 2024; He et al., 2024; Hsieh et al., 2024), Llama models might incorporate architectural modifications or training techniques designed to maintain robust attention across all positions in the context window, making them less susceptible to passage positioning issues.
Regarding distracting effects, when rotating a hard distractor among weak distracting passages, all models generally display the characteristic U-shaped pattern (see Figures 9, 10, 11 and 12), suggesting that distracting effects are more consistent across models and less influenced by parametric knowledge.
@@
LLM NQ TriviaQA PopQA
Q7B 44.20 68.80 20.40
L3B 58.60 68.00 20.60
L8B 67.40 80.80 30.60
L70B 74.60 92.20 49.60
B.2 Positional Bias in Real Scenarios
In Section 5, we presented experiments for $k$=5, showing minimal impact of different passage arrangement strategies on answer accuracy. Here, we expand the analysis to $k$=10 to investigate whether retrieving more passages might exhibit a more pronounced positional bias effect. Table 5 shows the answer accuracy across different LLMs when arranging the top-10 passages retrieved by BGE+RR using the strategies described in Section 5. Similar to the $k$=5 case, we observe that the positional bias has a marginal impact on answer accuracy. Across all LLMs, the difference between the best-performing strategy and the Shuffle strategy is not statistically significant according to the Wilcoxon test with $p$-value=0.05. Table 6 presents the results for Qwen 2.5 7B across different retrieval pipelines. We note one exception in the BGE retriever (without re-ranker), where the MaxRelevance strategy achieves 71.00 accuracy while Shuffle yields 68.87, which is a statistically significant difference. However, this appears to be an isolated case rather than a consistent pattern. This single exception does not contradict the broader statistical trend observed across all other configurations. For weaker retrievers like BM25, the positional ordering has less impact simply because they retrieve fewer relevant passages overall, as shown in Fig. 1a.
In general, these findings with $k$=10 reinforce our conclusion from the main paper: in realistic RAG settings, the impact of positional bias is minimal compared to its effect in controlled experimental conditions. The interaction between relevant and distracting passages in real retrieval results tends to neutralize potential benefits from strategic passage ordering.
@@ B.2 Positional Bias in Real Scenarios
{mdframed}
[font=]
You are given a question and you must respond based on the provided documents. Respond directly without providing any premise or explanation.Documents:<passage>...<passage>Question:<question>Answer:
@@
{mdframed}
[font=, innertopmargin=2pt, innerbottommargin=2pt, skipabove=0pt, skipbelow=0pt]

### v2

# Do RAG Systems Really Suffer From Positional Bias?

@@ Abstract
Retrieval Augmented Generation enhances LLM accuracy by adding passages retrieved from an external corpus to the LLM prompt. This paper investigates how positional bias–the tendency of LLMs to weight information differently based on its position in the prompt–affects not only the LLM’s capability to capitalize on relevant passages, but also its susceptibility to distracting passages. Through extensive experiments on three benchmarks, we show how state-of-the-art retrieval pipelines, while attempting to retrieve relevant passages, systematically bring highly distracting ones to the top ranks, with over 60% of queries containing at least one highly distracting passage among the top-10 retrieved passages. As a result, the impact of the LLM positional bias, which in controlled settings is often reported as very prominent by related works, is actually marginal in real scenarios since both relevant and distracting passages are, in turn, penalized. Indeed, our findings reveal that sophisticated strategies that attempt to rearrange the passages based on LLM positional preferences do not perform better than random shuffling.
Do RAG Systems Really Suffer From Positional Bias?
Florin Cuconasu1,2thanks: Work conducted while FC being a research intern at TII.thanks: cuconasu@diag.uniroma1.it, Simone Filice2 Guy Horowitz2, Yoelle Maarek2, Fabrizio Silvestri1 1Sapienza University of Rome, 2Technology Innovation Institute
@@ 1 Introduction
Retrieval Augmented Generation (RAG) improves the factual accuracy of LLMs on knowledge-intensive tasks by including in the prompt passages retrieved from an external corpus (Chen et al., 2017; Petroni et al., 2021b; Fan et al., 2024). Because any real retriever is imperfect, RAG systems feed the LLM several top-ranked passages, not just the single best one. That practice raises recall but also inserts distracting passages: text that looks relevant yet lacks the appropriate answer. Recent work shows these distractors can sharply degrade the LLM answer accuracy (Cuconasu et al., 2024; Jin et al., 2025; Yoran et al., 2024).
@@ 2 Related work
Recent work explores the detrimental effect of irrelevant content in the LLM prompt. In the RAG setting, a passage is considered irrelevant if it does not provide useful information for answering the query. Cuconasu et al. (2024) divide irrelevant passages as either random, if they are semantically unrelated to the query, or distracting, if they are related to the query but do not contain the answer. They show that while random passages do not affect answer quality, distracting passages do. Jin et al. (2025) show that irrelevant passages returned by strong retrievers are more detrimental than those obtained by weak retrievers. Amiraz et al. (2025) propose a continuous measure of the distracting effect of irrelevant passages and a fine-tuning approach to enhance LLM robustness, similar to strategies in Lin et al. (2024); Jin et al. (2025); Yoran et al. (2024). To mitigate these challenges, several approaches have emerged to compress or filter retrieved content: Yu et al. (2024) generate sequential reading notes that evaluate document relevance before answer generation, Xu et al. (2024) compress retrieved documents into concise textual summaries using both extractive and abstractive methods with selective augmentation, and Huang et al. (2024) perform sentence-level selection from encoded passages to reduce context length while preserving inference quality.
@@ 3 Experimental Setup
Following related work Zheng et al. (2023); Gu et al. (2025); Rahmani et al. (2024), we evaluate passage relevance and answer quality using the LLM-as-a-judge approach. In the former case, we prompt the LLM to assess the relevance of a passage to a question given the ground truth answer, where, following Cuconasu et al. (2024), we consider a passage relevant if and only if it contains the answer to the question, and irrelevant otherwise. In the latter, we prompt the LLM to assess whether the generated response semantically matches the reference answer [^3] Exact prompts are provided in Appendix A.3.. We use Claude 3.7 Sonnet via AWS Bedrock as the backbone LLM. [^4] We use Claude 3.7 Sonnet to minimize evaluation errors, though strong open-source models like Llama 3.3 70B Instruct achieve comparable performance (see Appendix A.3.1).
@@
$\text{DE}_{q}(p)=1-p^{\text{LLM}}(\text{NO-RESPONSE}|q,p)\vskip-2.84544pt$ (1)
For each retrieval pipeline, we compute the distracting effect only of irrelevant passages, and assign DE=0 for relevant passages. Fig. 1c reports the DE of the most distracting passage among the top-$k$ positions (MaxDE), while Fig. 1d reports the mean DE considering the top-$k$ positions (MeanDE). Both metrics are averaged across all queries. The MaxDE curves reach very high values, with Table 4 (Appendix) showing that over 60% of queries contain at least one hard distractor (defined as having a DE score greater than 0.8) in the top-10 results from dense retrievers. The MeanDE curves are initially very low, since most of the top retrieved passages are relevant, then increase as more irrelevant passages appear in the prompt, but soon they decrease again. This suggests that highly distracting passages typically appear in top positions, while low-ranked passages have a DE score close to 0. Finally, retrieval pipelines leading to higher HITS and Precision, e.g., when using BGE, also exhibit higher MaxDE and MeanDE curves, revealing a critical aspect: stronger retrievers increase recall and deliver more harmful distractors, making retrieval a double-edged sword.
You are given a question and you must respond based on the provided documents. Respond directly without providing any premise or explanation.
Documents:
DE=0.13 Document[1] (Title: Bids for the 2024 and 2028 Summer Olympics)(Section: Non-selected bids - 2024 - United States) [...] Following the final presentation, the USOC announced that the United States would bid to host the 2024 Olympic and Paralympic Games, but did not announce which city would bid. On 8 January 2015, the USOC selected Boston to be the candidate city from the United States but on 27 July 2015 Boston’s bid was withdrawn and the USOC bid process was reopened. On 1 September 2015 the USOC announced that Los Angeles was chosen for the United States bid for the 2024 Summer Games. DE=0.00 Document[2] (Title: Sports in the United States)(Section: Olympics) [...] The United States hosted both Summer and Winter Games in 1932, and has hosted more Games than any other country – eight times, four times each for the Summer and Winter Games: BULLET::::- the 1904 Summer Olympics in St. Louis, 1932 Summer Olympics and 1984 Summer Olympics in Los Angeles; and the 1996 Summer Olympics in Atlanta; BULLET::::- the 1932 Winter Olympics and 1980 Winter Olympics in Lake Placid, New York; the 1960 Winter Olympics in Squaw Valley, California; and the 2002 Winter Olympics in Salt Lake City, Utah. Los Angeles will host the Summer Olympics for a third time in 2028, marking the ninth time the U.S. hosts the Olympic Games. DE=0.01 Document[3] (Title: 1992 Winter Olympics)(Section: Legacy) The 1992 Olympic Winter Games marked the last time both the Winter and Summer games were held in the same year. The 1992 Olympics also marks the last time France hosted the Olympics. Paris will host the 2024 Summer Olympics. DE=0.19 Document[4] (Title: Sports in Chicago)(Section: Olympic bids) [...] Following Chicago’s loss in the race for the 2016 Olympics, the USOC bid for the 2024 Olympics with Los Angeles which result in a deal where Los Angeles secured the right to host the 2028 Summer Olympics. Chicago had previously hosted the 1959 Pan American Games. Chicago was selected to host the 1904 Summer Olympics, but they were transferred to St. Louis to coincide with the Louisiana Purchase Exposition. DE=0.98 Document[5] (Title: Summer Olympic Games)(Section: Hosting) The United States has hosted the Summer Olympic Games four times: the 1904 Games were held in St. Louis, Missouri; the 1932 and 1984 Games were both held in Los Angeles, California; and the 1996 Games were held in Atlanta, Georgia. The 2028 Games in Los Angeles will mark the fifth occasion on which the Summer Games have been hosted by the U.S. [...] Question: When did the united states host the last olympics?
Answer: The United States hosted the last Summer Olympics in 1996 in Atlanta, Georgia.
Gold Answer: 2002
Figure 3: Example showing how the position of the hard distractor affects Qwen 2.5 7B’s response when a relevant passage is fixed in position 2. When a hard distractor (Document 5, DE=0.98) is placed in position 5 (highest distracting effect according to Fig. 2b), the model provides an incorrect answer based on the hard distractor. However, simply moving the hard distractor to position 3 (lowest distracting effect), while maintaining the relevant passage in position 2, results in the model correctly answering “2002”.
@@ 4 Positional Bias in Controlled Settings
For each query, we select the highest-ranked relevant passage obtained by BGE after reranking (BGE+RR). Following Amiraz et al. (2025), we compute the distracting effect for irrelevant passages using Equation 1. We classify passages as “hard distractors” (with DE >0.8>0.8, as previously defined) and “weak distractors” (with DE <0.2<0.2). As an example, Fig. 3 illustrates this distinction: while weak distractors (Documents 1, 3, 4 with DE=0.01-0.19) have minimal impact on the model’s reasoning, hard distractors (Document 5 with DE=0.98) can cause the LLM to overlook correct information from relevant passages and generate incorrect answers. Fig. 2 shows results for Qwen 2.5 7B (results for other models and single datasets are given in Appendix B). Fig. 2a displays the characteristic U-shaped accuracy pattern when rotating a single relevant passage among fixed weak distractors [^5] We use weak distractors instead of general retrieved irrelevant passages to avoid negative effects from hard distractors.. Fig. 2b shows that this positional bias extends to distracting passages, with hard distractors at the beginning or end having significantly higher distracting effect (36-44%) compared to middle slots (28-34%) [^6] We calculate the distracting effect using Equation 1 applied to the entire set of passages rather than a single passage.. This parallel pattern indicates the model favors certain positions regardless of passage relevance.
@@ 5 Positional Bias in Real Scenarios
Results in Tables 2 and 3 show that the impact of the positional bias in real settings is minor: different passage arrangement strategies lead to very similar results that do not significantly differ from the Shuffle baseline [^9] Statistical significance using Wilcoxon test with pp=0.05., regardless of the LLM or the retrieval pipeline. We argue that these results can be explained by the contrastive effect of relevant and highly distracting passages, which, as observed in Fig. 1, tend to both appear in top retrieved passages: for instance, in the MaxRelevance strategy, the benefit of placing relevant passages in LLM-favored positions is compensated by the unintended tendency to put in the same slots highly distracting passages.
@@ 6 Conclusions
Our work demonstrates that while positional bias exists in current LLMs, its impact is minimal in realistic RAG settings: random ordering of retrieved passages yields statistically equivalent accuracy to more sophisticated reordering strategies. We observed that contemporary retrievers do not merely return some irrelevant passages, they surface passages that degrade answer accuracy in more than 60% of our test questions, turning the retriever itself into a first‑order source of error. Thus, attempting to place relevant passages in LLMs’ favorable positions may inadvertently prioritize hard distractors over relevant content, counterbalancing the potential benefits of strategic reordering. These findings suggest that future improvements should focus on retrieval quality and LLM distraction robustness rather than passage positioning.
@@
Acknowledgments
This work was carried out while Florin Cuconasu was enrolled in the Italian National Doctorate on Artificial Intelligence run by the Sapienza University of Rome. This project has also been supported by PNRR MUR project PE0000013-FAIR and PE00000014-SERICS, funded by European Union – NextGenerationEU, and the NEREO PRIN project funded by the Italian Ministry of Education and Research Grant no. 2022AEFHAZ.
@@ A.2 Additional Retrieval Results
Figures˜4, 5 and 6 report the retrieval results of BM25 and BGE with and without re-ranker (RR) on PopQA, NQ, and TriviaQA, respectively.
@@
A.2.1 Multi-Vector Retrieval with ColBERT
We also evaluated ColBERT (Khattab and Zaharia, 2020), a multi-vector retrieval method that performs fine-grained token-level matching between queries and passages. In terms of retrieval quality metrics (HITS@$k$, Precision@$k$, MaxDE@$k$, MeanDE@$k$), ColBERT exhibits trends similar to BGE (see Fig. 1). For the passage arrangement experiments in Section 5, ColBERT achieved the following accuracies with Qwen 2.5 7B on the top-5 retrieved passages: Sequential (67.13), Inverse (67.20), Shuffle (67.00), MaxRelevance (66.87), and MinDistraction (67.07). These results show no statistically significant differences between positional strategies (Wilcoxon test, pp=0.05), reinforcing our main conclusion that even sophisticated multi-vector retrievers surface a mixture of relevant and highly distracting passages, thereby mitigating positional bias effects in practical RAG scenarios.
A.3 LLM-as-a-Judge Methodology
A.3.1 Open-Source Alternative Validation
We also tested whether strong open-source models can serve as reliable alternatives to Claude 3.7 Sonnet for our evaluation tasks. We conducted a comparative analysis using Llama 3.3 70B Instruct on a subset of our data. We sampled 100 queries from each dataset (300 total) and evaluated relevance decisions for the top-5 passages retrieved by BGE using both models, analyzing a total of 1500 passages. For passage relevance assessment, the Cohen’s Kappa coefficient between Claude and Llama yielded a score of 0.85, indicating substantial agreement according to standard interpretation guidelines Landis and Koch (1977). For answer accuracy evaluation, we had both models assess the correctness of 300 generated answers. The agreement was even stronger, with a Cohen’s Kappa score of 0.94, indicating almost perfect agreement. These results suggest that Llama 3.3 70B Instruct can serve as a reliable open-source alternative for both passage relevance assessment and answer quality evaluation in RAG experiments.
@@ B.1 Positional Bias in Controlled Settings
Figures˜10, 11, 12 and 13 illustrate the positional bias in controlled settings when rotating either the relevant passage or a hard distractor among weak distractors. The results reveal that each model exhibits its own characteristic positional pattern, confirming findings from Hutter et al. (2025).
Among the LLMs tested, Qwen 2.5 7B demonstrates the most pronounced positional bias (see Fig. 10), while the Llama 3 family appears more resilient to position changes (see Figures˜11, 12 and 13). A possible explanation is that these models may have been specifically trained to mitigate the lost-in-the-middle effect. Since this problem has become well-documented in the literature (Liu et al., 2024; He et al., 2024; Hsieh et al., 2024), Llama models might incorporate architectural modifications or training techniques designed to maintain robust attention across all positions in the context window, making them less susceptible to passage positioning issues.
Regarding distracting effects, when rotating a hard distractor among weak distracting passages, all models generally display the characteristic U-shaped pattern (see Figures˜10, 11, 12 and 13), suggesting that distracting effects are more consistent across models and less influenced by parametric knowledge.
@@
B.2 Positional Bias in Real Scenarios
In Section 5, we presented experiments for kk=5, showing minimal impact of different passage arrangement strategies on answer accuracy. Here, we expand the analysis to kk=10 to investigate whether retrieving more passages might exhibit a more pronounced positional bias effect. Table 6 shows the answer accuracy across different LLMs when arranging the top-10 passages retrieved by BGE+RR using the strategies described in Section 5. Similar to the kk=5 case, we observe that the positional bias has a marginal impact on answer accuracy. Across all LLMs, the difference between the best-performing strategy and the Shuffle strategy is not statistically significant according to the Wilcoxon test with $p$-value=0.05. Table 7 presents the results for Qwen 2.5 7B across different retrieval pipelines. We note one exception in the BGE retriever (without re-ranker), where the MaxRelevance strategy achieves 71.00 accuracy while Shuffle yields 68.87, which is a statistically significant difference. However, this appears to be an isolated case rather than a consistent pattern. This single exception does not contradict the broader statistical trend observed across all other configurations. For weaker retrievers like BM25, the positional ordering has less impact simply because they retrieve fewer relevant passages overall, as shown in Fig. 1a.
In general, these findings with kk=10 reinforce our conclusion from the main paper: in realistic RAG settings, the impact of positional bias is minimal compared to its effect in controlled experimental conditions. The interaction between relevant and distracting passages in real retrieval results tends to neutralize potential benefits from strategic passage ordering.
@@ B.2 Positional Bias in Real Scenarios
{internallinenumbers*} You are given a question and you must respond based on the provided documents. Respond directly without providing any premise or explanation. Documents: <passage> ... <passage> Question: <question> Answer: